An object is placed in front of a convex mirror at distance of 50 cm. a plane mirror is introduced covering the lower half of the convex mirror. if the distance between the object and the plane mirror is 30 cm,it is found that there is no gap between the images formed by the 2 mirrors.the radius of the convex mirror is ?

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
Here the plane mirror and the convex mirror both form the image at the same position.
Object distance from the plane mirror=30 cm
So, Image distance =30 cm (Behind the mirror)
Thus formed image is at a distance of 10 cm behind the convex mirror.
Now,
For convex mirror,
u=-50 cm
v=10 cm
Using mirror formula,
1/v +1/u=1/f
1/10 -1/50=1/f
or f=12.5 cm
So R=2f=25 cm
 
Regards
Arun (askIITians forum expert)